To produce characteristic X – rays using a Tungsten target in an X-ray generator, the accelerating voltage should be greater than . . . . . . . . . . . . . . . . . . . .  volts and the energy of the characteristic radiation is . . . . . . . . eV.

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
For minimum accelerating voltage, the electron should jump from n = 2 to n = 1 level.
For characteristic X – rays
1 / λ = Rα (Z – 1 )2 [1 – 1 / n2] = = E / hc
∴ E1 / hc = Rα (Z - 1)2 [ 1 – 1 / 22] . . . . .(i)
The binding energy of innermost electron = 40 ke V
∴ Ionisation potential of tungsten = 40 k V = 40 x 103 V
⇒ E2 / E2 = [1 – 1/22] / [1 – 1/ ∞2]
⇒ E1 = ¾ E2 = ¾ x 40, 000 eV = 30, 000 eV
∴ Minimum accelerating voltage,
Vmin­ = E1 / e = 30, 000 V